Session: May 2004 (session 2)
Wavelength: 21 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Monari at Thu Jun 10 04:31:06 2004) Minor failures : HA-E
Everything OK during observations. Last source EL limited

Onsala (Michael Lindqvist at Thu Jun 10 04:51:43 2004) Minor failures :
Apart from RFI, no known problems. Source below horizon during the last three scans.

Jodrell_Bank (Alastair Gunn at Thu Jun 10 05:56:56 2004) Minor failures :
The Lovell Telescope was used. No known problems.

Westerbork (Tony Foley at Thu Jun 10 09:04:42 2004) Failure :
Wb used in phased array mode. Lost all observations after about 20:05UT. Airco problem meant that the IF and tied array racks stopped working. Sorry to all involbed.

Torun (K. Borkowski at Fri Jun 11 05:44:46 2004) Success :
No known problems.

Effelsberg (A. Kraus at Mon Jun 14 11:15:05 2004) Minor failures : WIND
Due to a thunderstorm the antenna was stopped between UT1811-2042.

Arecibo (Tapasi Ghosh at Tue Jun 15 18:47:32 2004) Success :
Not much troubles. However, for last 30 mins or so, the source had set for Arecibo, so the observation was ended by operators' intervention. This was to free the telescope for the next user.